One more then I'm done, my son has another problem on his homework that calls for the Gaussian elimination method with back substitution to solve the given system of these linear equations. would some be will to help explain this process a little clearer. the equations were -9x-45y= -27
x- y= -1
thanks in advance shannii

Answer by lynnlo(4164)

SUBSTITUTION METHOD:
-9x-45y=-27
x-y=-1
ANSWER:
y=2/3
x=-1/3